# The Unbounded Denominators Conjecture

### v1

@@ 4.2. Modular Forms
For an even integer $N$, we are considering the following rings of modular functions with rational coefficients and bounded denominators, that is, subrings of $\mathbf{Q}\otimes\mathbf{Z}\llbracket q^{1/N}\rrbracket$.
@@
    Let $M_{N}$ denote the ring of holomorphic modular functions on $Y(N)$ with coefficients in $\mathbf{Q}$.
@@
    Let $R_{N}$ denote the ring of holomorphic modular functions with coefficients in $\mathbf{Q}$, bounded denominators, and cusp widths dividing $N$.
Our goal is to prove that $R_{N}=M_{N}$. We can and do assume that $N$ is even. We have $M_{2}=\mathbf{Q}[\lambda^{\pm 1},(1-\lambda)^{\pm 1}]$ with function field $\mathbf{Q}(\lambda)$. The $\mathbf{Q}(\lambda)$-vector space of functions generated by elements of $R_{N}$ is finite dimensional. (We will later verify the assumptions of Proposition 3.0.1; note in any case that the crude finiteness $R_{N}:M_{2}]<\infty$ has already been proved from Theorem 2.0.2 and the remark that the conformal radius of $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is strictly larger than $1$.) It follows that all elements of $R_{N}$ are invariant under a subgroup $G_{N}$ of $\mathrm{SL}_{2}(\mathbf{Z})$ containing I-I which has finite index. In particular, $R_{N}$ is a subring of the holomorphic functions on the corresponding curve and so finitely generated over $M_{2}$. If $A$ is any finitely generated $M_{2}$-module, we write $A:M_{2}]=[A\otimes\mathbf{Q}(\lambda):\mathbf{Q}(\lambda)]$. Since $M_{2}$ has trivial Picard group, this is actually equivalent to $A$ being free of this rank (although we do not use this fact).
There are injective algebra maps $M_{2}\rightarrow M_{N}\rightarrow R_{N}$, where the degree of the first inclusion is given for $2|N$ by the explicit formula
@@
(4.2.2) $M_{N}:M_{2}]=\frac{1}{2}[\Gamma(2):\Gamma(N)]=\frac{N^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p|N}\left(1-\frac{1}{p^{2}}\right)>\frac{N^{3}}{12\zeta(2)}.$
Since modular forms on $Y(N)$ have bounded denominators, we furthermore have $R_{N}:M_{2}]=[R_{N}:M_{N}]\cdot[M_{N}:M_{2}]$. It follows that we have a bound:
@@
Combining the bounds (4.2.6) and (4.2.7) gives
$2^{(1-\varepsilon)X/\log X}<12C\zeta(2)X(1+\varepsilon)$
@@
We now put ourselves in the situation of Proposition 4.2.4. That is, we have $R_{N}$ is strictly larger than $M_{N}$. Recall that forms in $R_{N}$ are all invariant by a subgroup $G=G_{N}\subset\Gamma(N)$. By definition, all forms in $R_{N}$ have cusp width dividing $N$, and thus $G_{N}$ has level $N$ in the sense of Wohlfahrt. We may assume that $I\in G$. By Corollary 4.1.4, we may replace $G$ by a normal subgroup $G$ of finite index of $\mathrm{SL}_{2}(\mathbf{Z})$ containing I-I which also has finite index. Let $p$ be a prime with $(N,p)=1$.
The main idea of this section is to exploit the fact that if $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\in R_{N}$, then $f(\tau/p)\in\mathbf{Z}\llbracket q^{1/Np}\rrbracket$ is also a modular form with integer coefficients for any prime $p$. Since form $f(\tau)$ is invariant under $G$, the form $f(\tau/p)$ is invariant under $AGA^{-1}$ and thus also the group $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$. Now, by Lemma 4.1.7, we know that this group has (Wohlfahrt) level dividing $Np$. In particular $f(\tau/p)$ has cusp width dividing $Np$ at each cusp, and hence $f(\tau/p)\in R_{Np}$. Let $R_{N}M_{Np}$ denote the ring generated by forms in $R_{N}$ and $M_{Np}$, which is contained in $R_{Np}$. We have
@@
because the forms in $M_{Np}$ and $R_{N}$ lie in disjoint Galois extensions of (the field of fractions of) $M_{N}$. Thus
@@
is an integer, and it is either $\geq 2$, in which case we have proven Theorem 4.2.4, or $R_{Np}=R_{N}M_{Np}$. In particular, in order to prove Theorem 4.2.4 we may assume that $R_{Np}$ is generated by $R_{N}$ and $M_{Np}$, and in particular that $f(\tau/p)$ is invariant under the group $G\cap\Gamma(Np)$. This implies that $f(\tau/p)$ it is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. We will exploit an idea due to Serre that shows $G$ and $AGA^{-1}$ are in some sense “far apart” when $G$ is not a congruence subgroup. We first prove a preliminary lemma.
@@
Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a not-necessarily congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ with level $N$ in the sense of Definition 4.1.1. Let $H$ be the group generated by $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ together with $G\cap\Gamma(Np)$. Then $H$ contains $G\cap\Gamma(N)\cap\Gamma^{0}(p)$.
@@
There are maps
@@
The kernel of $\psi_{G}$ is equal to $G\cap\Gamma(Np)$, which is contained in $H$ by definition. Hence to show that $G\cap\Gamma(N)\cap\Gamma^{0}(p)$ is contained in $H$ it suffices to show that the image of $\psi_{H}$ contains the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$. Since $G$ has finite index inside $\mathrm{SL}_{2}(\mathbf{Z})$ and level $N$, it contains the matrices
$\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right),\quad\left(\begin{matrix}1&0\\ N&1\end{matrix}\right),$
and these generate the projection of $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ onto $\mathrm{SL}_{2}(\mathbf{F}_{p})$. Hence the image of $G\cap\Gamma(N)$ under $\psi_{G}$ is precisely the matrices $\gamma=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\mod Np$ congruent to $I\bmod N$. Now consider a matrix
@@
inside $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ which is congruent to $I\mod N$. These are precisely the matrices is in the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$, and our task is to show such matrices also lies the image of $H$. We first observe that the matrix
@@
also lies in $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ (the determinant is unchanged) and is congruent to the identity modulo $N$ since $(N,p)=1$. Thus it lies in the image of $G$. Let $\gamma$ denote any lift to $G$; clearly the lower left entry is divisible by $p$. Hence
@@
Since $G$ is normal, we may define the group $S$ by taking $S=\widetilde{\Gamma}(N)/G$. By construction, the group $S$ is finite. There are two homomorphisms $f_{1}$ and $f_{2}$ from $\widetilde{\Gamma}(N)\cap\Gamma_{0}(p)$ to $S$ defined by:
@@
    The inclusion map: $f_{1}:\widetilde{\Gamma}(N)\cap\Gamma_{0}(p)\rightarrow\widetilde{\Gamma}(N)\rightarrow S$,
@@
    The map $f_{2}=f_{1}(AxA^{-1})$.
Lemma 4.3.5 (Serre).
The map $(f_{1},f_{2}):\widetilde{\Gamma}(N)\cap\Gamma_{0}(p)\rightarrow S\times S$ is surjective.
@@
This follows as in the proof of [Tho89, Theorem 3] with the addition of $\widetilde{\Gamma}(N)$ level structure (See [Ber94] for the modification in the presence of level structure). The main point is that $(f_{1},f_{2})$ gives a map on the amalgamated product of $\widetilde{\Gamma}(N)$ and $A\widetilde{\Gamma}(N)A^{-1}$ on their intersection, and hence a map on the amalgamated product which may be identified with the level $N$ congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z}[1/p])$ which has the congruence subgroup property [Men67, Ser70]. ∎
@@
Let $K$ be the group generated by $\ker(f_{1})$ and $\ker(f_{2})$. Since $(f_{1},f_{2})$ is surjective, the image contains the elements $(x,0)$ and $(0,y)$ for any $x\in S$. But the pre-images of these elements clearly lie in $\ker(f_{2})$ and $\ker(f_{1})$ respectively, and thus lie in $K$. But then the pre-image of any element lies in $K$, and we deduce that
(4.3.6) $G\cap\Gamma_{0}(p)+A^{-1}GA\cap\Gamma_{0}(p)=\widetilde{\Gamma}(N)\cap\Gamma_{0}(p).$
We now consider the function $f(\tau/p)$, which is obtained from $f(\tau)$ by acting by $A^{-1}$. As explained at the beginning of § 4.3, we deduce that $f(\tau/p)$ is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\widetilde{\Gamma}(Np)$. But now conjugating by $A^{-1}$ (to bring $f(\tau/p)$ back to $f(\tau))$ we deduce from Lemma 4.3.1 that $f(\tau)$ is invariant under the group
@@
But it is also invariant under $G$, and hence tautologically under $G\cap\Gamma_{0}(p)$. By equation (4.3.6) it is thus also invariant under $\widetilde{\Gamma}(N)\cap\Gamma_{0}(p)$, which is to say invariant under a congruence subgroup, which was to be shown.

### v2

@@ 4.2. Modular Forms
For an even integer $N$, we will consider the following spaces of modular functions with rational coefficients generated by forms with bounded denominators, that is, subspaces of $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}=\mathbf{Q}\llbracket q^{1/N}\rrbracket[1/q]$ generated by elements of $\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes\mathbf{Q}$ as a $\mathbf{Q}(\lambda)$-vector space.
@@
    Let $M_{N}$ denote the $\mathbf{Q}(\lambda)$-vector space generated by holomorphic modular functions on the modular curve $Y(N)=\mathbf{H}/\langle E,\Gamma(N)\rangle$ with coefficients in $\mathbf{Q}$ at the cusp $\zeta=i\infty$.
@@
    Let $R_{N}$ denote the $\mathbf{Q}(\lambda)$-vector space generated by holomorphic modular functions with coefficients in $\mathbf{Q}$, bounded denominators at the cusp $\zeta=i\infty$, and cusp widths dividing $N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$.
(The vector space $R_{N}$ was also defined in Proposition 3.0.1 but we repeat the definition here for convenience.) For example, the holomorphic modular forms on $Y(2)$ are given by $\mathbf{Q}(\lambda^{\pm 1},(1-\lambda)^{\pm 1})$ and the $\mathbf{Q}(\lambda)$-vector space generated by such elements inside $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$ is $M_{2}=\mathbf{Q}(\lambda)$.
Lemma 4.2.2.
There is a containment $M_{N}\subset R_{N}$, and $M_{N}$ and $R_{N}$ have finite dimensions over $M_{2}=\mathbf{Q}(\lambda)$.
Proof.
Let $f$ be a holomorphic modular function on $Y(N)$, that is, a meromorphic function on the compact modular curve $X(N)$ whose poles are all at the cusps. Assume also that $f$ has coefficients in $\mathbf{Q}$ at the cusp $\zeta=i\infty$, Then the modular form $f\Delta(\tau)^{m}$ is holomorphic at the cusps for sufficiently large $m$. Moreover, $f\Delta(\tau)^{m}$ has coefficients in $\mathbf{Q}$. It follows from [Shi71, Theorem 3.52] that $f\Delta(\tau)^{m}$ has bounded denominators, and this implies the same for $f$, and thus there is a containment $M_{N}\subset R_{N}$.
The second claim follows from Corollary 2.0.5 and the remark that the conformal radius of $\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is strictly larger than $1$ given in the second paragraph of §1.1.3 (more precisely, $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is a well-defined holomorphic map with derivative being $1$ at 0 and thus the universal covering $D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has derivative strictly larger than $1$ by the Schwarz lemma). ∎
We have the following refinement of Lemma 4.2.2:
Lemma 4.2.3.
The vector spaces $M_{N}$ and $R_{N}$ are fields. The space $M_{N}$ may be identified with the field of rational functions on the modular curve $Y(N)/\mathbf{Q}$. There are injective algebra maps
$M_{2}\rightarrow M_{N}\rightarrow R_{N}.$
The space $R_{N}$ is invariant under a normal finite index subgroup $G_{N}\subset\langle E,\Gamma(N)\rangle\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(G_{N})=N$.
Proof.
Note that $M_{N}$ and $R_{N}$ are subspaces of $\mathbf{Q}(\kern-1.49994pt{(}q^{1/N})\kern-1.49994pt{)}$, which is a domain. Hence if $M_{N}$ and $R_{N}$ are rings then they are also integral domains, and any integral domain which has finite dimension over a field is itself a field.
The curve $Y(N)$ has a standard model over $\mathbf{Q}$ (as a moduli space of elliptic curves $C$ with a given symplectic isomorphism $C[N]\simeq\mathbf{Z}/N\mathbf{Z}\oplus\mu_{N}$) such that the cusp $\zeta=i\infty$ is defined over $\mathbf{Q}$, and the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$ on the global sections of $Y(N)$ is compatible with the $q$-expansion map. It follows that the set of generators of $M_{N}$ is closed under addition and multiplication and hence that $M_{N}$ is a ring, and thus a field. Moreover, $M_{N}$ contains the global sections of the (affine) curve $Y(N)/\mathbf{Q}$, and hence $M_{N}$ must be the function field of $Y(N)/\mathbf{Q}$.
The vector space $R_{N}$ is generated by holomorphic modular forms with bounded denominators at $\zeta=i\infty$. To show $R_{N}$ is a ring, it suffices to show that the product of any two such generators $g$ and $h$ is also a generator. Certainly $gh$ is a holomorphic modular form with rational coefficients and bounded denominators, so it suffices to show that the cusp width still divides $N$. But we may assume that $g$ and $h$ are invariant under finite index subgroup $G,H\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$, and thus $gh$ is invariant under $G\cap H$. It follows from Lemma 4.1.3 that $L(G\cap H)$ also has Wohlfahrt level dividing $N$.
Since $R_{N}$ is finite over $M_{2}$, it is generated by a finite number of basis elements each of which is invariant under some finite index subgroup $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ with $L(G)$ dividing $N$. The intersection of all these groups still has finite index and level $N$ by Lemma 4.1.3, and then we take $G_{N}$ to be the largest normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in this intersection, which also has $L(G_{N})=N$ by Lemma 4.1.4. ∎
Our goal is to prove that $R_{N}=M_{N}$. We can and do assume that $N$ is even. As noted in Lemma 4.2.3, $R_{N}:M_{2}]<\infty$.
The degree of $M_{N}$ over $M_{2}$ is equal to the degree of the modular curve $Y(N)$ over $Y(2)$, and this is given by the explicit formula
(4.2.4) $=\frac{1}{2}[\Gamma(2):\Gamma(N)]=$ $\ \frac{N^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p|N}\left(1-\frac{1}{p^{2}}\right)$
@@
$>$ $\ \frac{N^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p}\left(1-\frac{1}{p^{2}}\right)=\frac{N^{3}}{12\zeta(2)}.$
Since $R_{N}:M_{2}]=[R_{N}:M_{N}]\cdot[M_{N}:M_{2}]$, it follows that we have a bound:
@@
where the last inequality follows from the prime number theorem. Combining the bounds (4.2.8) and (4.2.9) gives
$2^{(1-\varepsilon)X/\log X}<12C\zeta(2)X(1+\varepsilon),$
@@
We now put ourselves in the situation of Theorem 4.2.6. That is, we have $R_{N}$ is strictly larger than $M_{N}$. Recall that forms in $R_{N}$ are all invariant by a subgroup $G=G_{N}\subset\langle E,\Gamma(N)\rangle$ with $L(G_{N})=N$ by Lemma 4.2.3. Let $p$ be a prime with $(N,p)=1$.
The main idea of this section is to exploit the fact that if $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\in R_{N}$, then $f(\tau/p)\in\mathbf{Z}\llbracket q^{1/Np}\rrbracket$ is also a modular form with integer coefficients for any prime $p$. Since the form $f(\tau)$ is invariant under $G$, the form $f(\tau/p)$ is invariant under $AGA^{-1}$ and thus also the group $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$. Now, by Lemma 4.1.7, we know that this group has (Wohlfahrt) level dividing $Np$. In particular $f(\tau/p)$ has cusp width dividing $Np$ at each cusp, and hence $f(\tau/p)\in R_{Np}$. Let $R_{N}M_{Np}$ denote the field generated by forms in $R_{N}$ and $M_{Np}$, which is contained in $R_{Np}$. We have
@@
because the intersection of $M_{Np}$ and $R_{N}$ is $M_{N}$. Thus
@@
is an integer, and it is either $\geq 2$, in which case we have proven Theorem 4.2.6, or $R_{Np}=R_{N}M_{Np}$. In particular, in order to prove Theorem 4.2.6, we may assume that $R_{Np}$ is generated by $R_{N}$ and $M_{Np}$, and in particular that $f(\tau/p)$ is invariant under the group $\langle E,G\cap\Gamma(Np)\rangle$. This implies that $f(\tau/p)$ is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. That is, Theorem 4.2.6 is reduced to the following claim:
Proposition 4.3.1.
If $f(\tau)$ is invariant by a finite index sugroup $G$ containing $E$ with $L(G)=N$ and $f(\tau/p)$ it is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$, then $f(\tau)$ is invariant under a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$.
The remainder of this section is devoted to the proof of this proposition. We will exploit an idea due to Serre [Tho89] that shows $G$ and $AGA^{-1}$ are in some sense “far apart” when $G$ is not a congruence subgroup.
We first prove a preliminary lemma. Recall that $\Gamma_{0}(p)$ and $\Gamma^{0}(p)$ are the subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$ consisting of matrices congruent to $\left(\begin{matrix}*&*\\ 0&*\end{matrix}\right)\bmod p$ and $\left(\begin{matrix}*&0\\ *&*\end{matrix}\right)\bmod p$ respectively. There is an identification $A\Gamma_{0}(p)A^{-1}=\Gamma^{0}(p)$ inside $\mathrm{SL}_{2}(\mathbf{Z})$.
@@
Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup with $L(G)=N$ in the sense of Definition 4.1.1. Let $H$ be the group generated by $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ together with $G\cap\Gamma(Np)$. Then $H$ contains $G\cap\Gamma(N)\cap\Gamma^{0}(p)$.
@@
Since $G$ and $H$ are both subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$, there are maps
@@
The kernel of $\psi_{G}$ is equal to $G\cap\Gamma(Np)$, which is contained in $H$ by definition. Hence to show that $G\cap\Gamma(N)\cap\Gamma^{0}(p)$ is contained in $H$, it suffices to show that the image of $\psi_{H}$ contains the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$. Since $G$ has finite index inside $\mathrm{SL}_{2}(\mathbf{Z})$ and Wohlfahrt level $N$, it contains the matrices
$\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right),\quad\left(\begin{matrix}1&0\\ N&1\end{matrix}\right).$
But it is easy to see that these two matrices generate the projection of $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ onto $\mathrm{SL}_{2}(\mathbf{F}_{p})$. Hence the image of $G\cap\Gamma(N)$ under $\psi_{G}$ is precisely the matrices $\gamma=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\mod Np$ congruent to $I\bmod N$, and the image of $G\cap\Gamma(N)\cap\Gamma^{0}(p)$ is precisely the matrices of the form
@@
inside $\mathrm{SL}_{2}(\mathbf{Z}/Np)$ which are congruent to $I\mod N$. Our task is to show that such matrices also lies the image of $H$. Since $(N,p)=1$, it follows that
(4.3.5) $b\equiv c\equiv 0\bmod N,\qquad a\equiv d\equiv 1\bmod N.$
We observe that the matrix
@@
also lies in $\mathrm{SL}_{2}(\mathbf{Z}/Np)$; the determinant is unchanged, and the matrix is congruent to the identity modulo $N$ by equation (4.3.5). Thus it lies in the image of $G\cap\Gamma(N)$. Let $\gamma$ denote any lift to $G$; clearly the lower left entry is divisible by $p$. Hence
@@
Since $G$ is normal and is contained in $\langle E,\Gamma(N)\rangle$, we may define the group $S$ by taking $S=\langle E,\Gamma(N)\rangle/G$. By construction, the group $S$ is finite. There is a natural projection:
$f:\langle E,\Gamma(N)\rangle\rightarrow\langle E,\Gamma(N)\rangle/G=S.$
We define two homomorphisms $f_{1}$ and $f_{2}$ from $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$ to $S$ as follows:
@@
    The map $f_{1}$ is the restriction of $f$ to $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$ under the natural inclusion
    $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow\langle E,\Gamma(N)\rangle,$
    so $f_{1}(x)=f(x)$.
@@
    Conjugation by $A$ induces an isomorphism
    $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow\langle E,\Gamma(N)\rangle\cap\Gamma^{0}(p),\quad\gamma\longrightarrow A\gamma A^{-1}.$
    The map $f_{2}$ is the composition of this map composed with $f$, so $f_{2}(x)=f(AxA^{-1})$.
Lemma 4.3.7 (Serre, Berger).
The map $(f_{1},f_{2}):\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow S\times S$ is surjective.
This is more or less precisely [Tho89, Theorem 3] with the addition of level structure as in [Ber94].
@@
The intersection of $\langle E,\Gamma(N)\rangle$ with $A\langle E,\Gamma(N)\rangle A^{-1}$ is the group $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$. We proceed by contradiction. Assume that the map $(f_{1},f_{2})$ is not surjective. By Goursat’s lemma, there exists a non-trivial quotient $\Delta$ of $S$ and projections $\pi_{i}:S\rightarrow\Delta$ such that the composites $\pi_{1}\circ f_{1}$ and $\pi_{2}\circ f_{2}$ agree. We define a map $g_{1}$ by the composite
(4.3.8) ${g_{1}:\langle E,\Gamma(N)\rangle\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{1}}$${\Delta}$
and a map $g_{2}$ by the composite
(4.3.9) ${g_{2}:A^{-1}\langle E,\Gamma(N)\rangle A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\langle E,\Gamma(N)\rangle\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{2}}$${\Delta}$
where the first map sends $x\rightarrow AxA^{-1}$. On the intersection
$\langle E,\Gamma(N)\rangle\cap A^{-1}\langle E,\Gamma(N)\rangle A=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p),$
the restriction of $g_{1}$ is given by $\pi_{1}\circ f_{1}$ and the restriction of $g_{2}$ is given by $\pi_{2}\circ f_{2}$. By construction these maps coincide, and hence they induce a surjective map on the amalgam
$\Phi:=\langle E,\Gamma(N)\rangle\star_{\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)}A^{-1}\langle E,\Gamma(N)\rangle A\rightarrow\Delta.$
There are natural inclusions from $\langle E,\Gamma(N)\rangle$ and $A^{-1}\langle E,\Gamma(N)\rangle A$ to the congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z}[1/p])$ consisting of matrices congruent to $\pm I\bmod N$, and these inclusions induce a map from $\Phi$ to this congruence subgroup. This map is an isomorphism ([Ber94, p.919], using ideas of [Ser80] and following the proof of [Tho89, Theorem 3]). But the group $\mathrm{SL}_{2}(\mathbf{Z}[1/p])$ (and thus the congruence subgroup $\Phi$) satisfies the congruence subgroup property [Men67, Ser70]. Hence the map $\Phi\rightarrow\Delta$ is a congruence map, and thus the same is true for the restriction to $\langle E,\Gamma(N)\rangle\subset\Phi$. This implies that the kernel $K\supseteq G$ of the map
$\langle E,\Gamma(N)\rangle\rightarrow\langle E,\Gamma(N)\rangle/G=S\rightarrow\Delta$
is a congruence subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ and strictly contained in $\langle E,\Gamma(N)\rangle$. But this contradicts the assumption that the Wohlfahrt level of $G$ is $N$, because the smallest congruence subgroup of Wohlfahrt level $N$ containing $E$ is precisely $\langle E,\Gamma(N)\rangle$ by [Woh64, Theorem 2]. ∎
@@
Let $K=\langle\ker(f_{1}),\ker(f_{2})\rangle$ be the group generated by $\ker(f_{1})$ and $\ker(f_{2})$. Since $(f_{1},f_{2})$ is surjective, the image of $(f_{1},f_{2})$ contains the elements $(x,0)$ and $(0,y)$ for any $x\in S$. But the pre-images of these elements clearly lie in $\ker(f_{2})$ and $\ker(f_{1})$ respectively, and thus lie in $K$. But then the pre-image of any element lies in $K$, and we deduce that $K=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, or equivalently that
(4.3.10) $\langle G\cap\Gamma_{0}(p),A^{-1}GA\cap\Gamma_{0}(p)\rangle=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p).$
We now complete the proof of Proposition 4.3.1 and hence the proof of Theorem 4.2.6 (as explained at the beginning of § 4.3).
Proof of Proposition 4.3.1.
Consider the function $f(\tau/p)$, which is invariant under both $AGA^{-1}\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. But now conjugating by $A^{-1}$ (to bring $f(\tau/p)$ back to $f(\tau))$ we deduce from Lemma 4.3.2 that $f(\tau)$ is invariant under the group
@@
But by assumption $f(\tau)$ is also invariant under $G$, and hence tautologically under $G\cap\Gamma_{0}(p)$. By equation (4.3.10), $f(\tau)$ is thus also invariant under $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, which is to say invariant under a congruence subgroup. ∎